# Can In-context Learning Really Generalize to Out-of-distribution Tasks?

### v1

@@ 3.2 Abstract Label Classification Can Only Be Achieved on In-Distribution Tasks
Figure 5: The ICL test error of Transformers evaluated on a quadratic regression + retrieval task. Different colors denote models trained on the linear regression + retrieval task with different numbers of label tokens. ’Eval’ denotes ’evaluated on’. The model trained on $s\sim\mathcal{U}(100,2000)$ doesn’t generalize better than the other two models.
3.3 Real-world LLMs Mat Not Necessarily In-context Learn New Tasks
Evaluating Llama2 on an OOD synthetic word classification task. In this section, we assess whether real-world LLMs can tackle OOD tasks through ICL. We select the pretrained Llama-2-7b and evaluate it on a synthetic word classification task. To ensure the task is far from the pretraining distribution, we randomly sample $\bm{x}_{i}\in\mathbb{R}^{d}$ from the word embedding of Llama-2-7b (denoted as $E_{llama}$) and generate random linear mappings $\bm{w}\in\mathbb{R}^{C\times d}$ as task functions (where $C=10$). The label words are created by mapping $\bm{x}_{i}$ to one of the ten label vectors in $E_{llama}$ using $\bm{w}$. Experimental details are in Appendix A.3. To complete this task, the model must learn $\bm{w}$ in context.
For comparison, we also evaluate the ICL performance of Llama-2-7b on a retrieval version of this task, where we sample $C$ classes of tokens from $E_{llama}$ to generate random sequences. The objective is to retrieve the same tokens as the query token from the context and output its subsequent token. The results of these two tasks are presented in Figure 6.
Observations. From Figure 6, we observe that the ICL performance on the synthetic classification task is close to random guessing (10% accuracy), while performance on the retrieval task is significantly better. Considering that the input and label distributions of the two tasks are very similar, we have reason to believe that LLMs struggle to learn new input-output mappings from context; instead, ICL appears to be more adept at retrieval tasks.
@@ 4 The Algorithm Selection Mechanism Exists Broadly When Evaluated on OOD Tasks
The Model pretrained on a single task vs. the model pretrained on multiple tasks. In Figure 7, we compare the performance of GPT-2 models trained on a single task—linear regression (LR), quadratic regression (QR), 2-layer ReLU network (ReLU NN) regression—against the model trained on all three tasks when encountering four kinds of OOD tasks. The results are in Figure 7.
Figure 7: The ICL performance of models trained on the individual task: linear regression (LR), quadratic regression (QR), 2-layer ReLU network (ReLU NN) regression, and the model trained on the mixture of the three tasks (LR+QR+ReLU NN). The evaluation functions are (a) square root, (b) cubic, (c) linear+quadratic, and (d) 2-layer Sigmoid network (details in Appendix A.1). The performance of the model trained on the mixed tasks is comparable to that of the model trained on the single task that performs the best on the evaluation task.
Observations. From Figure 7, we observe that the ICL performance of the model trained on mixed tasks is comparable to the performance of the model trained on a single task with the lowest test error on the evaluation task. ICL seems to automatically select the best pretraining functions according to the downstream context.
4.1 Theoretical Analysis
A mixed Gaussian pretraining dataset of multiple tasks. In this section, we theoretically analyze the algorithm selection mechanism of ICL on OOD tasks, based on the theoretical framework of Lin & Lee (2024). Consider a noisy linear regression pretraining dataset with the inputs and task weights following the mixed Gaussian distribution:
@@ 4.1 Theoretical Analysis
The lemma below states that the closed-form prediction of the model trained on the pretraining data under Assumption 4.1, given the testing context, remains a Gaussian mixture of the reweighted pretraining task weights.
Lemma 4.2.
(Corollary 2. of Lin & Lee (2024), closed-form ICL prediction of the pretrained model) Denote the model $M^{*}$ that minimizes the risk on the pretraining data of Assumption 4.1, i.e., $M^{*}\in\arg\min\frac{1}{T}\sum_{i=0}^{T-1}\underset{S_{i}\sim P_{tr}}{\mathbb{E}}\left[\left\|M\left(\mathcal{S}_{i}\oplus\bm{x}_{i+1}\right)-y_{i+1}\right\|^{2}\right]$, then the prediction on any sequence $\mathcal{S}_{i}\oplus\bm{x}_{i+1}$ by $M^{*}$ is as follows: $M^{*}\left(\mathcal{S}_{i}\oplus\bm{x}_{i+1}\right)=\left\langle\bm{x}_{i+1},\sum_{m=1}^{M}\tilde{\pi}_{m}\tilde{\bm{w}}_{m}\right\rangle$. where $\tilde{\pi}_{m}$, and $\tilde{\bm{w}}_{m}$ depending on both the pretraining task and the downstream context example are given in Lemma 1 of Lin & Lee (2024).
Based on the closed-form ICL prediction, we now analyze how the downstream context affects $\tilde{\pi}$, which determines how ICL selects the pretraining functions. First, we introduce Lemma 4.3 that characterizes the ratio of the reweighted weight of two pretraining tasks:
@@
(ICL prediction favors the pretraining function with low error on the context) Given the context $\bm{S}_{T}$, if the empirical risk of implementing a function of the pretraining task $\alpha$ is less than that of $\beta$, i.e., $\frac{1}{T}\sum_{i=1}^{T}|\bm{w}_{\beta}\bm{x}_{i}-y_{i}|^{2}-|\bm{w}_{\alpha}\bm{x}_{i}-y_{i}|^{2}\geq 0$, then, under some mild Assumptions C.2 on the distribution of $\bm{S}_{T}$, we have $\Psi_{\bm{w}}(\alpha,\beta)\geq 0$.
@@
Summary of the algorithm-selection mechanism. 4.3 and Theorem 4.4 together elucidate the algorithm-selection mechanism of ICL. According to Lemma 4.2, the ICL prediction of the model pretrained on the mixed Gaussian data will be a reweighted combination of the pretraining task vectors $\bm{w}_{i}$. Whether the ratio between the weights of two pretraining tasks, $\tilde{\pi}_{\alpha}/\tilde{\pi}_{\beta}$, given a downstream context, exceeds the original ratio $\pi_{\alpha}/\pi_{\beta}$ depends on two factors: 1) whether the pretraining input distribution of $\alpha$ is closer to the downstream input distribution than that of $\beta$; 2) whether the task function of $\alpha$ induces lower test error in downstream context than that of $\beta$. When both conditions are met, we have $\tilde{\pi}_{\alpha}/\tilde{\pi}_{\beta}\geq\pi_{\alpha}/\pi_{\beta}$, indicating that ICL prefers $\alpha$ over $\beta$ in its predictions.
Discussion of the setup of our theory. Notably, our theoretical result doesn’t assume a Transformer architecture, while previous theoretical works of understanding ICL often adopt Transformers with oversimplified assumptions on their parameters or structures (Ahn et al., 2024; Zhang et al., 2023; Huang et al., 2023; Collins et al., 2024). Additionally, our analysis shows that models pretrained on the ICL tasks can implement algorithm selection during ICL inference following Lin & Lee (2024). In contrast, prior work on algorithm selection (Bai et al., 2024) only shows that a specific set of parameters in a simplified ReLU Transformer can enable algorithm selection. However, the parameter construction is complex and somewhat tricky, and there is no theoretical or experimental guarantee that Transformers exhibiting algorithm selection will necessarily implement these parameters.
@@ 4.2 Empirical Validation of the Algorithm-Selection Mechanism of ICL
Now we validate our theoretical findings regarding ICL’s algorithm-selection mechanism in OOD tasks by conducting numerical experiments following Lin & Lee (2024). In Figure 8 and 9, the training data is a Gaussian mixture with four components (see Assumption 4.1), while the test function is a two-layer ReLU network (Appendix A.1). Both the training and the test data are in ICL format. We compute the test error of using each pretraining task function to predict the downstream function (the first row of Figure 8 and 9), the weights for each pretraining function during ICL inference (the second row), and the test error of the pretrained ICL model (the third row). We evaluate five different noise levels ($\delta_{x}=\delta_{w}\in\{1/81,9/1,1,9,81\}$) and consider two settings described below.
Low-test-error preference of ICL. To validate Theorem 4.4, we ensure that the distributional distances between the inputs of each training task and the test data remain consistent. Specifically, all $\bm{x}_{i}$ in both training data and test data are sampled from $\mathcal{N}([0,0,0]^{\top},\sigma^{2}_{x}\bm{I})$. The task weights for different pretraining tasks vary, as detailed in the top half of Table 1. In this setup, only the test error of the pretraining functions influences algorithm selection. From the top two rows of Figure 8, we can observe a clear negative correlation between the ICL performance and the test error of the task weight. This result supports Theorem 4.4, confirming that ICL prefers the pretraining functions with a low test error in the downstream context. Also, it’s consistent with the observations in Figure 7.
Table 1: Experiment setting of Figure 8 and Figure 9. ’PT’ and ’DS’ are short for ’pretraining’ and ’downstream’, respectively.
@@
Figure 8 $\mathcal{N}([0,0,0]^{\top},\sigma^{2}_{x}\bm{I})$ 1 $\mathcal{N}([0,0,0]^{\top},\sigma^{2}_{x}\bm{I})$ $\mathcal{N}([5,5,5]^{\top},\sigma^{2}_{w}\bm{I})$ 0
@@
Figure 9 $\mathcal{N}([-4,-4,-4]^{\top},\sigma^{2}_{x}\bm{I})$ 1 $\mathcal{N}([5,5,5]^{\top},\sigma^{2}_{w}\bm{I})$ $\mathcal{N}([1,1,1]^{\top},\sigma^{2}_{w}\bm{I})$ 15.59
@@
Figure 8: Empirical validation of the low-test-error preference of ICL. The test error of the four pretraining functions in this experiment varies (see the first row), while the distributional distances between the inputs of each training task and the test data remain consistent. In the first two rows, a clear negative correlation between the test error of pretraining task $i$ and the task weight $\tilde{\pi}_{i}$ can be observed.
Similar-input-distribution preference of ICL. We also empirically validate Lemma 4.3 in Figure 9. In this case, the distributional distances between the input of different pretraining tasks and that of the test context vary, while the test errors of different pretraining functions are almost the same (detailed setup is in the bottom half of Table 1). As shown in the bottom two rows in Figure 9, the task weight $\tilde{\pi}_{i}$ is positively correlated with the similarity between the training and testing input distribution. This is consistent with Lemma 4.3 which demonstrates that ICL prefers to select the pretraining function whose input distribution is close to the downstream one.
Figure 9: Empirical validation of the similar-input-distribution preference of ICL. The test error of the four pretraining functions remains almost consistent (see the first row), while the distributional distances between the inputs of the four training tasks and the test data vary. The distances of different tasks are ordered from largest to smallest as $1>2>3>4$. From the first two rows, we observe that the task weight $\tilde{\pi}_{i}$ is negatively correlated with the distance between the training and testing input distribution.
5 Comparison with Related Works
Abilities to learn new tasks of ICL. Besides studies indicating that ICL can learn new weights of linear regression (Garg et al., 2022; Raventós et al., 2024; Zhang et al., 2023; Akyürek et al., 2022), other research has found that LLMs can tackle tasks that are unlikely to have been encountered during pretraining. For example, Pan (2023) showed that LLMs perform better than random guessing on classification tasks with meaningless labels. Kossen et al. (2024) demonstrated that ICL can identify authorship based on writing style in private communication messages not included in the pretraining corpus. Additionally, Vacareanu et al. (2024) found that large-scale LLMs can learn various linear and non-linear functions from context. We argue that these findings do not contradict our work. While the LLMs may not have seen exactly the same tasks, there is no guarantee that they haven’t encountered tasks from a similar distribution in their pretraining corpus. For instance, the LLMs could have been pretrained on a corpus containing authorship identification tasks or on statistical data encompassing different functions. Our work does not claim that ICL cannot generalize to new task instances; rather, it highlights the limitation in generalizing to an unseen input-label distribution. Additionally, Yadlowsky et al. (2023) finds that ICL struggles to generalize to testing function classes that are unseen during training (e.g., convex combinations or extreme versions of the pretraining functions). They didn’t delve into how ICL behaves on OOD data, while we reveal that it implements the pretraining functions.
The algorithm-selection mechanism of ICL. Recent works by Bai et al. (2024); Wang et al. (2024) have uncovered the algorithm selection phenomenon, demonstrating that Transformers pretrained on both linear regression and classification tasks perform well when presented with the context of either task during ICL inference. Theoretically, they showed that a Transformer with specific parameters can achieve algorithm selection. Yadlowsky et al. (2023) empirically found that ICL selects the optimal pretraining function class after observing in-context examples from a function class present in the pretraining data mixture. However, the algorithm selection experiments in these studies are limited to scenarios where the test functions are among the training functions. In this work, we empirically and theoretically demonstrate that the algorithm selection phenomenon broadly occurs when given downstream context from arbitrary function classes. To the best of our knowledge, we are the first to reveal the factors that determine the selection process.

### v3

@@ 4.2 Abstract Label Classification Can Only Be Achieved on ID Tasks
Figure 5: The ICL test error of Transformers evaluated on a quadratic regression + retrieval task. Different colors denote models trained on the linear regression + retrieval task with different numbers of label tokens. “Eval” denotes “evaluation”. The model trained on $s\sim\mathcal{U}(100,2000)$ doesn’t generalize better than the other two models.
Summary of the Empirical Results & Connections with the Existing Theories (II) To handle classification with abstract labels, the model infers an input-label mapping to implicitly establish a similarity metric. It then retrieves in-context examples similar to the query to deduce the OOD labels using this metric. However, this process succeeds only when the underlying function is ID, thus it does not represent a true OOD generalization capability. This observation aligns with Bayesian frameworks for ICL—the implicit similarity metric here corresponds to the task concept inferred by the model. We leave an intuitive Bayesian interpretation of the findings in Section 4.1 and 4.2 in Appendix A.4.
4.3 Real-world LLMs May Not Necessarily In-context Learn New Tasks
Evaluating Llama-3 on an OOD synthetic word classification task. Now we assess whether real-world LLMs can tackle OOD tasks through ICL. We design a synthetic word classification task for a pretrained Llama-3-8B. Specifically, we randomly sample $\bm{x}_{i}\in\mathbb{R}^{d}$ from the word embedding of Llama-3-8B (denoted as $E_{llama}$) and generate random linear mappings $\bm{W}\in\mathbb{R}^{d\times C}$ as task functions (where $C=10$). The label words are created by mapping $\bm{x}_{i}$ to one of the ten label vectors in $E_{llama}$ using $\bm{W}$. Experimental details are in Appendix C.4. To complete this task, the model must learn $\bm{W}$ in context, which is unlikely to have been seen during the pretraining of Llama.
For comparison, we also evaluate the ICL performance of Llama-3-8B on a retrieval version of this task. Concretely, we first randomly sample $C=10$ different vectors from $E_{llama}$ as $\bm{x}_{i}$ and compute $\bm{y}_{i}$ in the same way as the above classification task to get $S=[\bm{x}_{1},\bm{y}_{1},...,\bm{x}_{C},\bm{y}_{C}]$. Then we repeat $S$ 20 times to construct the input sequence $S^{\prime}=[S\oplus S\oplus...\oplus S]$, where $\oplus$ denotes concatenation operation. The goal is to predict the next token given a prefix of $S^{\prime}$. To succeed in this task, the model has to retrieve the same token as the query token (the last $\bm{x}_{i}$ of $S^{\prime}$) from the context and output its subsequent token $\bm{y}_{i}$. The results of these two tasks are presented in Figure 6.
Observations. From Figure 6, we observe that the ICL performance on the synthetic classification task is close to random guessing (10% accuracy), while performance on the retrieval task is significantly better (similar results also hold for Llama-2-7B in Appendix B.3). This suggests that pretrained real-world LLMs may also struggle to learn new input-output mappings from context; instead, ICL appears to be more adept at retrieval tasks. To show that the failure in the synthetic word classification task is mainly due to its OOD nature instead of some other factors that make it difficult to learn, we train a GPT-2 to perform the same task in Appendix B.2 and find that the task can be well addressed after training.
Figure 6: The ICL accuracy of Llama-3-8B on the synthetic tasks.
@@ 5 The Algorithm Selection Mechanism Exists Broadly When Evaluated on OOD Tasks
The Model pretrained on a single task vs. the model pretrained on multiple tasks. In Figure 7, we compare the performance of GPT-2 models trained on a single task—linear regression (LR), quadratic regression (QR), 2-layer ReLU network (ReLU NN) regression—against the model trained on all three tasks when encountering four kinds of OOD tasks. We also plot the error of a 2-layer ReLU NN trained by GD (dashed blue line). The results are in Figure 7.
Figure 7: The ICL performance of models trained on the individual task: linear regression (LR), quadratic regression (QR), 2-layer ReLU network (ReLU NN) regression, and the model trained on the mixture of the three tasks (LR+QR+ReLU NN). The evaluation functions are (a) square root, (b) cubic, (c) linear+quadratic, and (d) 2-layer Sigmoid network (details in Appendix C.1). The performance of the model trained on the mixed tasks is comparable to that of the model trained on the single task that performs the best on the evaluation task.
Observations. 1) the ICL performance of the model trained on mixed tasks (LR+QR+ReLU NN) is comparable to the performance of the model trained on a single task with the lowest test error on the evaluation task. This suggests that ICL can automatically select the best pretraining functions according to the downstream context. 2) ReLU NN consistently performs the best on all four OOD test functions. Moreover, the performance of the ReLU model trained by GD aligns well with the ICL performance of the GPT-2 trained on the same function class. This demonstrates that our findings in Section 3.1 still hold when the Transformer is trained on a mixture of multiple tasks.
5.1 Theoretically Revealing the Mechanism of Algorithm Selection
In this section, we will provide theoretical insight into the working mechanism of the algorithm selection of ICL. We find there simultaneously exist two parallel mechanisms: the Low-test-error preference and the Similar-input-distribution preference.
A mixed Gaussian pretraining dataset of multiple tasks. We adopt the theoretical framework by Lin & Lee (2024). Consider a noisy linear regression pretraining dataset with the inputs and task weights following the mixed Gaussian distribution:
@@ 5.1 Theoretically Revealing the Mechanism of Algorithm Selection
According to the Corollary 2 of Lin & Lee (2024) (see E.1 in Appendix E), the closed-form prediction of the model trained on the pretraining data under Assumption 5.1, given the testing context, remains a Gaussian mixture of the reweighted pretraining task weights: $\left\langle\bm{x}_{i+1},\sum_{m=1}^{M}\tilde{\pi}_{m}\tilde{\bm{w}}_{m}\right\rangle$, where $\tilde{\pi}_{m}$ and $\tilde{\bm{w}}_{m}$ are the posterior variables of $\pi_{m}$ and $\bm{w}_{m}$ given the downstream context. Hence, to analyze how ICL selects pretraining priors, the key lies in uncovering how $\tilde{\pi}_{m}$ evolves. First, we introduce Lemma 5.2 that characterizes the ratio of the reweighted weight of two pretraining tasks:
@@
(ICL prediction favors the pretraining function with low error on the context) Given the context $\bm{S}_{T}$, if the empirical risk of implementing a function of the pretraining task $\alpha$ is less than that of $\beta$, i.e., $\frac{1}{T}\sum_{i=1}^{T}|\bm{w}_{\beta}\bm{x}_{i}-y_{i}|^{2}-|\bm{w}_{\alpha}\bm{x}_{i}-y_{i}|^{2}\geq 0$, then, under some mild Assumptions E.2 on the distribution of $\bm{S}_{T}$, we have $\Psi_{\bm{w}}(\alpha,\beta)\geq 0$.
@@
Summary of the algorithm-selection mechanism. 5.2 and Theorem 5.3 together elucidate the algorithm-selection mechanism of ICL. According to Lemma E.1, the ICL prediction of the model pretrained on the mixed Gaussian data will be a reweighted combination of the pretraining task vectors $\bm{w}_{i}$. Whether the ratio between the weights of two pretraining tasks, $\tilde{\pi}_{\alpha}/\tilde{\pi}_{\beta}$, given a downstream context, exceeds the original ratio $\pi_{\alpha}/\pi_{\beta}$ depends on two factors: 1) whether the pretraining input distribution of $\alpha$ is closer to the downstream input distribution than that of $\beta$; 2) whether the task function of $\alpha$ induces lower test error in downstream context than that of $\beta$. When both conditions are met, we have $\tilde{\pi}_{\alpha}/\tilde{\pi}_{\beta}\geq\pi_{\alpha}/\pi_{\beta}$, indicating that ICL prefers $\alpha$ over $\beta$ in its predictions. We leave the discussions of the advantage of our theory result in Appendix A.5 and offer an intuitive Bayesian interpretation of the algorithm selection in Appendix A.4.
@@ 5.2 Empirical Validation of the Algorithm-selection Mechanism of ICL
Now we validate our theoretical findings regarding ICL’s algorithm-selection mechanism in OOD tasks by conducting numerical experiments following Lin & Lee (2024). In Figure 8(a) and 8(b), the training data is a Gaussian mixture with four components (see Assumption 5.1), while the test function is a two-layer ReLU network (Appendix C.1). Both the training and the test data are in ICL format. We compute the test error of using each pretraining task function to predict the downstream function (the first row of Figure 8(a) and 8(b)), the weights for each pretraining function during ICL inference (the second row), and the test error of the pretrained ICL model with the closed form prediction derived in Lemma E.1 (the third row). We evaluate five different noise levels ($\delta_{x}=\delta_{w}\in\{1/81,9/1,1,9,81\}$) and consider two settings described below.
Low-test-error preference of ICL. To validate Theorem 5.3, we ensure that the distributional distances between the inputs of each training task and the test data remain consistent. Specifically, all $\bm{x}_{i}$ in both training and test data are sampled from $\mathcal{N}([0,0,0]^{\top},\sigma^{2}_{x}\bm{I})$. The task weights for different pretraining tasks vary, as detailed in the top half of Table 1. In this setup, only the test error of the pretraining functions influences algorithm selection. From the top two rows of Figure 8(a), we can observe a clear negative correlation between the ICL performance and the test error of the task weight. This result supports Theorem 5.3, confirming that ICL prefers the pretraining functions with a low test error in the downstream context. Also, it’s consistent with the observations in Figure 7.
Similar-input-distribution preference of ICL. We also empirically validate Lemma 5.2 in Figure 8(b). In this case, the distributional distances between the input of different pretraining tasks and that of the test context vary: the distances of different tasks are ordered from largest to smallest as $1>2>3>4$, while the test errors of different pretraining functions are almost the same (detailed setup is in the bottom half of Table 1). As shown in the bottom two rows in Figure 8(b), the task weight $\tilde{\pi}_{i}$ is positively correlated with the similarity between the training and testing input distribution. This is consistent with Lemma 5.2 which demonstrates that ICL prefers to select the pretraining function whose input distribution is close to the downstream one.
(a) Numerical verification of the low-test-error preference
(b) Numerical verification of the similar-input-distribution preference
Figure 8: Empirical validation of the algorithm-selection mechanism of ICL. The first rows: the test error of the four pretraining functions. The mid rows: the weights of each pretraining function in the closed-form downstream ICL prediction (given by Lemma E.1). The last rows: the test error of the pretrained ICL model with the closed form prediction derived in Lemma E.1. Observations. 1) In the first two rows of Figure 8(a), the value of the task weight $\tilde{\pi}_{i}$ is negatively correlated with the test error of pretraining task $i$. 2) In the first two rows of Figure 8(b) the task weights are negatively correlated with the distance between the training and testing input distribution.
Table 1: Experiment setting of Figure 8(a) and Figure 8(b). “PT” and “DS” are short for “pretraining” and “downstream”, respectively.
@@
Figure 8(a) $\mathcal{N}([0,0,0]^{\top},\sigma^{2}_{x}\bm{I})$ 1 $\mathcal{N}([0,0,0]^{\top},\sigma^{2}_{x}\bm{I})$ $\mathcal{N}([5,5,5]^{\top},\sigma^{2}_{w}\bm{I})$ 0
@@
Figure 8(b) $\mathcal{N}([-4,-4,-4]^{\top},\sigma^{2}_{x}\bm{I})$ 1 $\mathcal{N}([5,5,5]^{\top},\sigma^{2}_{w}\bm{I})$ $\mathcal{N}([1,1,1]^{\top},\sigma^{2}_{w}\bm{I})$ 15.59
@@
5.3 Verifying the Algorithm-selection Mechanism on Real-world LLMs
In this section, we investigate whether real-world LLMs can perform algorithm selection through ICL. To achieve this, we design an ambiguous sentence classification task, in which each sentence can be classified based on one of three aspects: “sentiment”, “type”, or “location”. For each ICL sequence, we select one of the aspects as the classification criterion and map the label words to meaningless strings. For instance, if we choose to classify each sentence according to its sentiment, then “positive,” “neutral,” and “negative” are mapped to “RqF,” “IwZ,” and “SdK,” respectively. Detailed experimental setups are in Appendix C.5. We compute the top-5 accuracy of different classification criteria. The results in Figure 9 show that as the context length increases, the LLM finds the most appropriate criterion, exhibiting the low-test-error preference.
(a) Sentiment
(b) Type
(c) Location
Figure 9: The top-5 accuracy of using (a)“sentiment”, (b)“type”, or (c)“location” as the classification criterion for in-context examples in a test prompt. The accuracy of using the true underlying criterion to predict is significantly higher than the other two. This suggests that LLMs can perform algorithm selection in natural language tasks.
Summary of the Empirical Results & Connections with the Existing Theories (III) Despite the impressive empirical performance of real-world LLMs in solving seemingly novel tasks through ICL, we observe that when faced with an OOD task, ICL operates by identifying the most suitable pretraining meta-distribution based on test error and input distribution discrepancies, and then attempts to find an optimal solution within that meta-distribution. Notably, this process occurs independently of the downstream test distribution. See Appendix A.4 for the potential connection between such empirical observations and the Bayesian framework work.